The P/E ratio of the S&P 500 index is 20. The dividend payout ratio of the S&P 500 index is 30%. The expected annual growth rate of dividends for the portfolio of S&P 500 stocks is 2%. What is your estimated expected return from investing in the S&P 500?
Answer #1

P/E ratio = 20/1

dividend payout ratio = 30%

growth = 2%

RRR or expected return =(Expected dividend payment/Share price) ​+Forecasted dividend growth rate

= (30/20) + 0.02

= 1.52 %